Macroeconomic

Macroeconomic

media%2Fb69%2Fb69360db-5e3b-41e9-8bbc-e7
Answer #1

RGDP in 2016 = 14461 billion

RGDP in 2017 = 14702 billion

Population in 2016 = 121 million = 0.121 billion

Population in 2017 = 122 million = 0.122 billion

a) Growth rate of RGDP =( RGDP (2017/2016) -1)*100

Growth rate of RGDP =( (14702/14461)-1)x100

Growth rate of RGDP = 1.67%

B) GDP per person = Real GDP / Population

GDP per person in 2016 = 14461/0.121 = $119512

GDP per person in 2017 = 14702/0.122 = $120508

Growth rate = ((120508/119512)-1)*100

Growth rate = 0.833%

C) Rule 70 is used to find GDP growth rate

Number of years to double RGDP = 70/ %Growth rate of RGDP

N = 70/1.67

N = 41.92 years.